OpenStudy (anonymous):

Kim's coffee shop sells a cup of tea for $1.05, a cup of coffee for $1.35, and a cup of hot chocolate for $2.25 . One busy day, 20 more cups of coffee than of hot chocolate were sold, and 30 more cups of coffee than tea, for a total of 202.50 for all three hot drinks. How many cups of each drink were sold?

OpenStudy (amistre64):

T=1.05; C=1.35; Hc=2.25. C = Hc+20 ; C=T+30. Hc + 20 = T+30 sooo... Hc = T +30-20 Hc = T + 10 this is what I got so far :)

OpenStudy (amistre64):

(T+30)(1.35) + (T+10)(2.25) + T(1.05) = 202.50 If I did it right :)

OpenStudy (amistre64):

1.35T + 40.5 + 2.25T +22.5 + 1.05T = 202.50

OpenStudy (amistre64):

T(1.35 +2.25+1.05) +63 = 202.50 T(4.65) = 202.50 -63 T = 139.5/4.65 T = 30. There were 30 teas sold.

OpenStudy (amistre64):

C=T+30 and Hc = T+10 60 coffees sold; and 40 hot chocolates sold.

OpenStudy (amistre64):

Am I right?

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yur a genius

OpenStudy (amistre64):

awww.... you only saying that cause its true :)

OpenStudy (anonymous):

yah :) thnks again

OpenStudy (anonymous):

kso , I'm doing the same question , but I don't get how you got " 1.35T + 40.5 + 2.25T +22.5 + 1.05T = 202.50" ?